markel7:

Points A and B lie on a circle centered at point O. If OA = 5 and what is the area of sector AOB? Use the value π = 3.14, and choose the closest answer.

AZ:

we're going to need more information- probably missing a diagram which tells us the angle \(\theta\) but this is all that I can surmise from the information given  \(\text{Area} = \pi r^2 \dfrac{\theta}{360}\)

AZ:

we know that r = 5 but we need more information, possibly an image you forgot to attach, to figure out what the angle is and then we can find the area of the sector AOB but hopefully the formula lets you get started on the right track @Markel7